If there are 2.50 moles of an unknown atomic element and the mass of this amount is 60.78 g, the unknown element must be:


A - Sodium


B - Magnesium


C - Aluminium


D - Silicon

The mass of a sample of 0.124 mol of tin(II) fluoride is _____g. Remember to provide your answer using significant digits.


In 50g of ammonia, there are _____ mol. Remember to use significant digits for your answer


The molar mass of iron (III) sulfate is ______ g/mol.


The expected products after combining sodium iodide with lead(II) nitrate would likely be:


A - NaNO3(aq) + PbI(s)


B - Na(NO3)2(aq) + PbI2(s)


C - Na(NO3)2(aq) + PbI(s)


D - NaNO3(aq) + PbI2(s)


Convert 540000000 to scientific notation in the form of a.b X 10c. Enter the digits a, b, and c in this order for your 3-digit numerical answer.


When performing the calculation below, the final answer would be rounded to 10, according to the rounding and significant digit rules.


   (4.36 X 2.2) + 0.8580


True or false?




1
Expert's answer
2021-11-16T15:39:03-0500

Solution:

(1):

The molar mass of he unknown element = (60.78 g) / (2.50 mol) = 24.312 g/mol

It is magnesium (Mg)

Answer (1): B - Magnesium


(2):

tin(II) fluoride = SnF2

The molar mass of SnF2 is 156.69 g/mol

Hence,

(0.124 mol) × (156.69 g mol-1) = 19.42956 g SnF2 = 19.4 g SnF2

Answer (2): 19.4 g


(3):

ammonia = NH3

The molar mass of NH3 is 17.031 g/mol

Hence,

(50 g NH3) × (1 mol NH3 / 17.031 g NH3) = 2.9356 mol NH3 = 3 mol NH3

Answer (3): 3 mol


(4):

iron(III) sulfate = Fe2(SO4)3

The molar mass of iron(III) sulfate is 399.88 g/mol

Answer (4): 399.88 g/mol


(5):

sodium iodide = NaI

lead(II) nitrate = Pb(NO3)2

Balanced chemical equation: 2NaI(aq) + Pb(NO3)2(aq) → 2NaNO3(aq) + PbI2(s)

Answer (5): D - NaNO3(aq) + PbI2(s)


(6):

540000000 = 5.4×108

a.b×10c

(a = 5; b = 4; c = 8) = 548

Answer (6): 548


(7):

(4.36 × 2.2) + 0.8580 = 10.45 = 10.5 - True

Answer (7): True